# How Does Critical Batch Size Scale in Pre-training?

### v2

@@ 1.2 Theoretical Implications
Consider mini-batch SGD with $D$ samples in the least square problems under power-law source and capacity conditions. The CBS, which enables mini-batch SGD to achieve the minimal expected excess risk while ensuring the fastest possible serial runtime, is given by $B^{*}(D)=\Theta(D^{c})$, where the exponent $c\geq 0$ is determined by the exponents of the source and capacity conditions. In the regime where the variance error tends to be dominant, we have $c>0$, indicating CBS grows with data size.
@@ 2 Experimental Design and Empirical Findings
We first describe the experimental settings and refer readers to Appendix D for extra details. Throughout this paper, we use the abbreviations ‘M’ for million, ‘B’ for billion, and ‘T’ for trillion.
@@ 2.1 Experimental Settings
Experimental design and outline. To study CBS in Chinchilla settings, we need to consider target loss on a holdout validation set and measure the number of optimization steps required to reach it. We consider the optimal batch size as $B_{\text{opt}}$ that is in the linear scaling regime that incurs no efficiency overhead as detailed in Appendix C and model architectures in Table 2. We consider the validation loss of an optimal batch size $B_{\text{opt}}=256$ at step $t_{\textrm{Chin}}=C_{\textrm{Chin}}*N/(\textrm{ctx\_len}*B)$ for each model size $N$ and batch size $B$ with context length ctx_len set to be 512.
@@ 2.2 Training Beyond Fixed Durations for Reaching Target Validation Loss
Benchmarking learning rate schedulers. In practice, LMs are usually trained using fixed token budgets (Hoffmann et al., 2022a), which can determine the total number of iterations the training would undergo. This training process can be easily decomposed into learning rate warmup and decay phases so that a lower learning rate is kept at the end of training to enable better optimization. However, our goal is to find the optimal performing run under various hyper-parameters and optimization conditions. This implies a non-trivial decision regarding selecting the maximum training duration (Defazio et al., 2024; Hägele et al., 2024). As training beyond fixed durations is particularly favorable in many large-scale pre-training scenarios, we benchmark recently proposed methods like schedule-free optimizer (Defazio et al., 2024), cosine, warmup-stable-decay schedule (WSD) (Hu et al., 2024) (or trapezoidal (Zhai et al., 2022)) and our proposed constant+EWA strategy which maintains a running average of model weights $\xi_{t+1}=\tau\cdot\xi_{t}+(1-\tau)\cdot\theta_{t}$ to improve optimization and use $\xi$ for evaluations. We show that our constant+EWA strategy can match the efficiency of cosine scheduling and WSD, especially for large batch sizes (Figure 2).
@@
Takeaway on learning rate scheduling: EWA consistently improves model training efficiency. Training with EWA is slightly worse than Cosine learning rate decay for small batch sizes while outperforming Cosine for large batch sizes ($>$0.52M tokens) Cosine scheduling, while beneficial for small-batch training, proves sub-optimal for large-batch scenarios. Its reliance on a fixed-duration schedule limits its adaptability, and even with appropriate learning rate decay, it underperforms our constant+EWA strategy in large-batch settings.
@@ 2.4 Ablation on Model Width and Depth
Figure 4: Scaling up width and depth are equivalent in efficiency gain for compute-optimal training.
Firstly, as shown in Figure 4 (left, middle), under Chinchilla settings where data and model size are scaled proportionally, increasing either model depth or width leads to a similar increase in the CBS. Then through controlled comparison (Figure 4, right), we see that using two different ways to scale 151M models to 604M ones is equivalent in efficiency since both curves overlap.
Our findings may offer practical insights for scaling models under a fixed token budget that is allocated in proportion to model size. This is particularly relevant because scaling model width is often favored over increasing depth, as wider models tend to be more amenable to parallelization without incurring additional latency overhead (Shoeybi et al., 2019; Touvron et al., 2023; Erdil, 2024).
Takeaway on scaling transformer width and depth in compute-optimal regimes: Increasing width and depth has similar effects in the increase of critical batch size for compute-optimal pre-training.
@@ 3 Critical Batch Size Scaling Law
3.1 Formal Definition of Critical Batch Size
Recall that CBS is the transition point where increasing the batch size by a factor of $k$, leads to a reduction in the required number of training steps by a factor that is less than $k$. We now define CBS as the batch size that leads to a $20\%$ overhead compared to linear scaling. First of all, define $\mathcal{R}(N,D,B)$ as the best loss achievable for a model of size $N$ using a single pass on $D$ tokens with a batch size $B$. This would be obtained by optimally tuning all other parameters of the optimizer, while keeping $N,D,B$ fixed. Below is the formal definition of CBS:
@@ 5 Related Work
Previous studies have shown that increasing batch sizes can be offset by a proportional adjustment to the learning rate in small-scale regimes (McCandlish et al., 2018; Zhang et al., 2019; Kaplan et al., 2020). McCandlish et al. (2018) introduce the gradient noise scale, a measure that captures the variation in gradients across different training examples, which helps predict the critical batch size (CBS). Their findings also suggest that small-batch training is more compute-efficient, while large-batch training requires fewer optimizer steps. Momentum-based methods extend scaling to larger batch sizes but converge to the performance of standard SGD at smaller batch sizes (Shallue et al., 2019). Additionally, Zhang et al. (2019) analyze the impact of curvature on CBS using a noisy quadratic model, demonstrating that preconditioning techniques can increase the CBS. Golmant et al. (2018) show that the size of the dataset plays a smaller role in determining training efficiency compared to factors like model architecture and data complexity. In contrast, Hilton et al. (2022) examine how performance can be maintained at smaller batch sizes. Meanwhile, Smith et al. (2017); Smith and Le (2017) empirically investigated how the optimal learning rate changes based on momentum and training set size. Theoretical work has further sought to characterize CBS by analyzing SGD behavior in least-squares linear regression, especially in over-parameterized settings (Jain et al., 2018; Ma et al., 2018). Filatov et al. (2024) concurrently find that optimal batch size and CBS scale with data size. However, they do not explore how CBS scales with model size for models beyond 354M parameters, nor do they provide theoretical justifications or address the challenge of selecting optimal runs across a broad range of hyperparameters. Our work advances the optimization literature by formalizing CBS and quantifying its growth w.r.t. data size and emphasizing the importance of common hyper-parameter choices. It also provides strategies for studying large-scale pre-training beyond fixed training durations.
@@ 6 Concluding Remarks
In conclusion, this study provides an extensive examination of the scaling laws for critical batch size in large-scale autoregressive language model pre-training. By systematically analyzing the relationship between model size, data size, and CBS, we found that while CBS increases with data size, it remains relatively invariant to model size. This finding suggests training on more data may enable greater data parallelism in pre-training. We further emphasize the role of key hyperparameters and exponential weight averaging, which can match the performance cosine scheduling without requiring fixed training durations. These insights offer practical strategies for scaling models while maintaining efficiency, which is critical in resource-constrained scenarios.
@@ Appendix A Training Dynamics
Scheduler comparison for other batch sizes. In Figure 9, we include more comparisons on different schedulers that are reported in the main text (Figure 1). Overall, our Constant+EWA performs competitively with cosine scheduling and outperforms WSD scheduling, especially for large batch size regimes. Note that we sweep over the decay steps as 0.1,0.2,0.3×0.1,0.2,0.3$\times$0.1 , 0.2 , 0.3 × total training steps for WSD scheduling. We tune cosine scheduling by conducting sweeps over various maximum optimization steps to identify the optimal value, and then rerun the training using this step count. This approach ensures that the model reaches the target loss near the end of training, optimizing the performance of learning rate decay. Under small batch sizes, the schedule-free optimizer (Defazio et al., 2024) is a competitive baseline but it is significantly worse for batch sizes larger than 1024.
@@
(h) Batch size 8192
Figure 9: Scheduler comparison for all the smaller batch sizes. All are with model size 151M.
@@ Appendix B Additional Ablation Studies on Studying Adam Optimizer
We employ Adam as the default optimizer for large-scale model training throughout the paper. In this section, we focus on two key hyper-parameters that significantly affect optimization efficiency and examine their impact in detail.
@@ Appendix C Results Including Small Batch Sizes
Moreover, we include all the results that contain the smallest several batch sizes (Figure 13). Note that the denominator is the number of steps to reach target loss at batch size 64 instead of 256 now. Now we can observe clear linear scaling of all the model sizes till around $2^{10}$ for model sizes, while the largest three model sizes maintain linear scaling till almost $2^{11}$. There are minor differences with the main plot in Figure 1 because of the difficulty of optimizing with very small batch sizes like 64 but it does not affect the conclusions and takeaways we would like to convey. Since our focus is primarily on large batch sizes, we consistently use $2^{9}$ as the starting batch size throughout the main text.
Furthermore, recall that we set $B_{\text{opt}}=2^{8}$ when selecting the target loss. Figure 12 and Figure 13 confirm that $2^{8}$ falls within the linear scaling regime, which justifies our design choice.
@@ Appendix D Additional Details on Experiments
Optimizer setup. For optimizers, we try both SGD (Robbins and Monro, 1951) and Adam (Kingma, 2014) and find that SGD without momentum is significantly worse so we use Adam only for all the experiments. We disable weight decay in Adam. For generality, though the training set C4 might contain low-quality or duplicated documents that can potentially lead to training instability (Muennighoff et al., 2023; Wortsman et al., 2023), we observed that these issues did not affect our primary target of interest—namely, the final optimization efficiency. As a result, we didn’t explicitly adopt additional normalization like QK normalization (Dehghani et al., 2023; Zhai et al., 2023) or a z-loss (Chowdhery et al., 2022) to mitigate loss spikes. [^2] We observe irregular loss spikes despite adopting gradient clipping, but most runs can still be optimized well in the end. We set ϵ=italic-$\epsilon=$1.0e-8 by default, and refer to momentum as $\beta_{1}$ in Adam by default throughout the paper.
@@
Evaluation data size and frequency. As we need to frequently evaluate the model on a holdout evaluation set, understanding how large evaluation set size can lead to a stable and reliable measure yet still pertain efficiency of each run can be important. We evaluated the variance across different token counts: 2.17e-4 for 327,680 tokens, 4.53e-5 for 1,638,400 tokens, and 7.65e-6 for 3,276,800 tokens. Based on these results, we have set the default number of evaluation batches to 100.
It is important to note that the total number of training steps varies significantly across different batch sizes. To address this, we implement a hybrid evaluation protocol: the model is evaluated at intervals of $2^{i}$ (where $i\in\mathbb{Z}$ ), every 1,000 steps, and at $0.7n,0.75n,0.8n,\ldots$, up to $n$ during the last $30\%$ of the total steps $n$. This approach ensures more frequent evaluations toward the end of training, allowing for a more accurate assessment of the total training steps needed to achieve a target evaluation loss.
@@
Table 3: Sweeping experiments settings. Default values after the hyper-parameter search are in Bold font. Values in the parathesis mean used not for every sweep.
@@ Appendix E Additional Details on Scaling Laws
We first present the fitted power law relationship between the number of optimization steps required to reach the target loss and the batch size (Table 6)
@@
Note that our definition of CBS and its scaling law have a similar interpretation with the one in (McCandlish et al., 2018; Kaplan et al., 2020) as $\frac{E_{\min}}{S_{\min}}$, ${S_{\min}}$ denotes the minimum possible number of steps taken to reach target loss and $E_{\min}$ is the minimum possible number of training examples processed to reach target loss. In particular, recall that critical batch size can be analytically derived as $B^{*}=\frac{b}{5a}+1.2B_{\mathrm{opt}}$. This relationship reflects the point where batch size scaling incurs a 20% overhead when the batch size is doubled while (McCandlish et al., 2018). Here, the parameter $b$ plays a role analogous to $E_{\min}$, while $a$ corresponds to $S_{\min}$, depending on the specific overhead chosen to characterize the diminishing returns from increasing the batch size.

### v3

@@ 1.2 Theoretical Implications
Consider mini-batch SGD with $D$ samples in the least square problems under power-law source and capacity conditions. The CBS, which enables mini-batch SGD to achieve the minimal expected excess risk while ensuring the fastest possible serial runtime, is given by $B^{*}(D)=\Theta(D^{c})$, where the exponent $c\geq 0$ is determined by the exponents of the source and capacity conditions. In the regime where the variance error tends to be dominant, we have $20<c<1/$ 2, indicating CBS grows with data size.
@@ 2 Experimental Design and Empirical Findings
We describe the experimental settings and refer readers to Appendix D for extra details. Throughout this paper, we use the abbreviations ‘M’ for million, ‘B’ for billion, and ‘T’ for trillion.
@@ 2.1 Experimental Settings
Experimental design and outline. To study CBS in Chinchilla settings, we need to consider target loss on a holdout validation set and measure the number of optimization steps required to reach it. We consider the optimal batch size as $B_{\text{opt}}$ in the linear scaling regime that incurs no efficiency overhead as detailed in Appendix C. We consider the validation loss of an optimal batch size $B_{\text{opt}}=256$ at step $t_{\textrm{Chin}}=C_{\textrm{Chin}}\times N/(\textrm{ctx\_len}\times B)$ for each model size $N$ and batch size $B$ with context length ctx_len set to be 512, where $C_{\textrm{Chin}}$ is the Chinchilla coefficient.
@@ 2.2 Training Beyond Fixed Durations for Reaching Target Validation Loss
Benchmarking learning rate schedulers. In practice, LMs are usually trained using fixed token budgets (Hoffmann et al., 2022a), which can determine the total number of iterations the training would undergo. This training process can be easily decomposed into learning rate warmup and decay phases so that a lower learning rate is kept at the end of training to enable better optimization. However, our goal is to find the optimal performing run under various hyper-parameters and optimization conditions. This implies a non-trivial decision regarding selecting the maximum training duration (Defazio et al., 2024; Hägele et al., 2024). As training beyond fixed durations is particularly favorable in many large-scale pre-training scenarios, we benchmark recently proposed methods like schedule-free optimizer (Defazio et al., 2024), cosine, warmup-stable-decay schedule (WSD) (Hu et al., 2024) (or trapezoidal (Zhai et al., 2022)) and our proposed constant+EWA strategy which maintains a running average of model weights $\xi_{t+1}=\tau\cdot\xi_{t}+(1-\tau)\cdot\theta_{t}$ to improve optimization, where $\theta_{t}$ is the actual model parameter at step $t$ and we use $\xi$ for evaluation. To ensure the baselines are close to optimal, we first get the optimal step counts from our constant+EWA runs, we evaluate WSD scheduling by testing [0.1, 0.2, 0.3] multiples of those step counts. In parallel, for the cosine scheduler, we explore [0.9, 1.0, 1.1] times the same step counts. We show that our constant+EWA strategy can match the efficiency of cosine scheduling and WSD, especially for large batch sizes (Figure 2). The connections are explained in detail in (Morwani et al., 2025).
@@
Takeaway on learning rate scheduling: EWA consistently improves model training efficiency compared to using a constant learning rate without it. EWA proves to be an effective approach compared to other baselines with decaying schemes, offering competitive performance while eliminating the need to predefine training durations.
@@ 2.4 Ablation on Model Width and Depth
Figure 3: Scaling up width and depth shares similar efficiency gain for compute-optimal training.
@@
Firstly, as shown in Figure 3 (left, middle), under Chinchilla scaling—where data and model size grow proportionally— increasing model depth or width has a similar impact on CBS. Notably, according to previous results, the rise in CBS in compute-optimal scaling should be attributed to the increased data size. Then through controlled comparison (Figure 3, right), we see that using two different ways to scale 151M models to 604M ones is equivalent in efficiency since both curves overlap. Our findings may offer practical insights for scaling models under a fixed token budget that is allocated in proportion to model size. This is particularly relevant because scaling model width is often favored over increasing depth, as wider models tend to be more amenable to parallelization without incurring additional latency overhead (Shoeybi et al., 2019; Touvron et al., 2023; Erdil, 2024; McLeish et al., 2025).
Takeaway on scaling transformer width and depth in compute-optimal regimes: Increasing width and depth has similar effects in critical batch size for compute-optimal pre-training.
@@ 3 Critical Batch Size Scaling Law
3.1 Formal Definition of Critical Batch Size
Recall that CBS is the transition point where increasing the batch size by a factor of $k$, leads to a reduction in the required number of training steps by a factor that is less than $k$. We now define CBS as the batch size that leads to a $20\%$ overhead compared to linear scaling.
@@ 3.1 Formal Definition of Critical Batch Size
First of all, define $\mathcal{R}(N,D,B)$ as the best loss achievable for a model of size $N$ using a single pass on $D$ tokens with a batch size $B$. This would be obtained by optimally tuning all other parameters of the optimizer, while keeping $N,D,B$ fixed. Below is the formal definition of CBS:
@@ 5 Related Work
Previous studies have shown that increasing batch sizes can be offset by a proportional adjustment to the learning rate in small-scale regimes (McCandlish et al., 2018; Zhang et al., 2019; Kaplan et al., 2020; Li et al., 2021). McCandlish et al. (2018) introduce the gradient noise scale, a measure that captures the variation in gradients across different training examples, which helps predict the critical batch size (CBS). Their findings also suggest that small-batch training is more compute-efficient, while large-batch training requires fewer optimizer steps. Momentum-based methods extend scaling to larger batch sizes but converge to the performance of standard SGD at smaller batch sizes (Shallue et al., 2019). Additionally, Zhang et al. (2019) analyze the impact of curvature on CBS using a noisy quadratic model, demonstrating that preconditioning techniques can increase the CBS. Golmant et al. (2018) show that the size of the dataset plays a smaller role in determining training efficiency compared to factors like model architecture and data complexity. In contrast, Hilton et al. (2022) examine how performance can be maintained at smaller batch sizes. Meanwhile, Smith et al. (2017); Smith and Le (2017) empirically investigated how the optimal learning rate changes based on momentum and training set size. Theoretical work has further sought to characterize CBS by analyzing SGD behavior in least-squares linear regression, especially in over-parameterized settings (Jain et al., 2018; Ma et al., 2018). Filatov et al. (2024) concurrently find that optimal batch size and CBS scale with data size. However, they do not explore how CBS scales with model size for models beyond 354M parameters, nor do they provide theoretical justifications or address the challenge of selecting optimal runs across a broad range of hyperparameters. Our work advances the optimization literature by formalizing CBS and quantifying its growth w.r.t. data size and emphasizing the importance of common hyper-parameter choices. It also provides strategies for studying large-scale pre-training beyond fixed training durations.
@@ 6 Concluding Remarks
In conclusion, this study provides an extensive examination of the scaling laws for critical batch size in large-scale autoregressive language model pre-training. By systematically analyzing the relationship between model size, data size, and CBS, we found that while CBS increases with data size, it remains relatively invariant to model size. This finding suggests training on more data may enable greater data parallelism in pre-training. We further emphasize the role of key hyperparameters and exponential weight averaging, which can match the performance of cosine scheduling without requiring fixed training durations. These insights offer practical strategies for scaling models while maintaining efficiency, which is critical in resource-constrained scenarios.
@@ Appendix A Training Dynamics
Scheduler comparison for other batch sizes. In Figure 9, we include more comparisons on different schedulers that are reported in the main text (Figure 1). Overall, our Constant+EWA performs competitively with cosine scheduling and outperforms WSD scheduling, especially for large batch size regimes. Note that we sweep over the decay steps as 0.1,0.2,0.3×0.1,0.2,0.3$\times$0.1 , 0.2 , 0.3 × total training steps for WSD scheduling. We tune cosine scheduling by conducting sweeps over various maximum optimization steps to identify the optimal value, and then rerun the training using this step count. This approach ensures that the model reaches the target loss near the end of training, optimizing the performance of learning rate decay. For schedule-free optimizers, we tune the $\beta_{1}$ 0.9, 0.95, 0.98. Under small batch sizes, the schedule-free optimizer (Defazio et al., 2024) is a competitive baseline but it is significantly worse for batch sizes larger than 1024.
@@
Figure 9: Scheduler comparison for different batch sizes. All are with model size 151M.
@@
Effects of weight decay. Though weight decay (WD) does not provide generalization benefits for pre-training, previous works have shown that it might improve convergence in language model training (Hoffmann et al., 2022b; Kosson et al., 2023). We adopt the default decoupled weight decay (Loshchilov, 2017) implementation in PyTorch and sweep over weight decay rate $\{0.01,0.0316,0.1\}$ for LR $0.01$, $\{0.0316,0.1,0.316\}$ for LR $0.00316$. We show in Figure 11 that for constant learning rate with EWA, while weight decay offers a slight performance improvement, it has little effect on the critical batch size, which remains our primary focus. Consequently, we disable weight decay throughout the paper.
@@ Appendix B Additional Ablation Studies on Studying Adam Optimizer
We employ Adam as the default optimizer for large-scale model training throughout the paper.
Figure 11: Comparison of efficiency with and without weight decay
In this section, we focus on two key hyper-parameters that significantly affect optimization efficiency and examine their impact in detail.
@@ Appendix C Results Including Small Batch Sizes
Moreover, we include all the results that contain the smallest several batch sizes (Figure 14). Note that the denominator is the number of steps to reach target loss at batch size 64 instead of 256 now. Now we can observe clear linear scaling of all the model sizes till around $2^{10}$ for model sizes, while the largest three model sizes maintain linear scaling till almost $2^{11}$. There are minor differences with the main plot in Figure 1 because of the difficulty of optimizing with very small batch sizes like 64 but it does not affect the conclusions and takeaways we would like to convey. Since our focus is primarily on large batch sizes, we consistently use $2^{9}$ as the starting batch size throughout the main text. Furthermore, recall that we set $B_{\text{opt}}=2^{8}$ when selecting the target loss. Figure 13 and Figure 14 confirm that $2^{8}$ falls within the linear scaling regime, which justifies our design choice.
@@ Appendix D Additional Details on Experiments
Optimizer setup. For optimizers, we try both SGD (Robbins and Monro, 1951) and Adam (Kingma, 2014) and find that SGD without momentum is significantly worse so we use Adam only for all the experiments. We disable weight decay in Adam as we observe that it does not significantly affect critical batch size (Figure 11). For generality, though the training set C4 might contain low-quality or duplicated documents that can potentially lead to training instability (Muennighoff et al., 2023; Wortsman et al., 2023), we observed that these issues did not affect our primary target of interest—namely, the final optimization efficiency. As a result, we didn’t explicitly adopt additional normalization like QK normalization (Dehghani et al., 2023; Zhai et al., 2023) or a z-loss (Chowdhery et al., 2022) to mitigate loss spikes. [^2] We observe irregular loss spikes despite adopting gradient clipping, but most runs can still be optimized well in the end. We set $\epsilon$ to be 1e-8 by default, and refer to momentum as $\beta_{1}$ in Adam by default throughout the paper.
@@
Evaluation data size and frequency. To ensure frequent model evaluation on a holdout C4 validation set, it is important to maintain a balance between reliability and efficiency. A larger evaluation set size can provide more stable and reliable performance metrics, but it must also be efficient to maintain practicality in each run. Using 151M models, we evaluated the variance across different token counts: 2.17e-4 for 327,680 tokens, 4.53e-5 for 1,638,400 tokens, and 7.65e-6 for 3,276,800 tokens. Based on these results, we have set the default number of evaluation batches to 100. It is important to note that the total number of training steps varies across different batch sizes. To address this, we implement a hybrid evaluation protocol: the model is evaluated at intervals of $2^{i}$ (where $i\in\mathbb{Z}$ ), every 1,000 steps, and at $0.7n,0.75n,0.8n,\ldots$, up to $n$ during the last $30\%$ of the total steps $n$. This approach ensures more frequent evaluations toward the end of training, allowing for a more accurate assessment of the total training steps needed to achieve a target evaluation loss.
@@
Table 3: Sweeping experiments settings. Default values after the hyper-parameter search are in Bold font. Bold font means the default hyper-parameters that can closely reproduce our results without extensive tuning. Not bolding implies a full sweep for each model scale. The values in parentheses were not used for every sweep: for the 151M models, we tested learning rates of 3.16e-4 and 1e-2, but found that with EWA, these performed worse than 3.16e-3. The EWA decay rate of 0.99995 is only used for long 1.2B runs.
@@ Appendix E Additional Details on Scaling Laws
We first present the fitted power law relationship between the number of optimization steps required to reach the target loss and the batch size (Table 6). All the results are obtained by solving the equation in Section 3.2 via scipy.optimize.fsolve using default hyper-parameters.
@@
Note that our definition of CBS and its scaling law have a similar interpretation with the one in (McCandlish et al., 2018; Kaplan et al., 2020) as $\frac{E_{\min}}{S_{\min}}$, ${S_{\min}}$ denotes the minimum possible number of steps taken to reach target loss and $E_{\min}$ is the minimum possible number of training examples processed to reach target loss. In particular, recall that critical batch size can be analytically derived as $B^{*}=\frac{b}{5a}+1.2B_{\mathrm{opt}}$. This relationship reflects the point where batch size scaling incurs a 20% overhead when the batch size is doubled while (McCandlish et al., 2018). Here, the parameter $b$ plays a role analogous to $E_{\min}$, while $a$ corresponds to $S_{\min}$, depending on the specific overhead chosen to characterize the diminishing returns from increasing the batch size. We also note that the diminishing return overhead can vary, leading to the following observations: $10\%:B^{*}=20.67\times D^{0.48},20\%:B^{*}=22.91\times D^{0.47},50\%:B^{*}=30.50\times D^{0.44}$.